OpenStudy (anonymous):

How are the sin 20° and the cos 70° related? Explain using complete sentences.

OpenStudy (anonymous):

@Ryan2416 please help me with this on tooo. (:

OpenStudy (jdoe0001):

http://ichthyosapiens.com/School/Math/Trig/ch05CofunctionIdentities.jpg <--- see the relationship?

OpenStudy (jdoe0001):

\(\bf cos(90^o-\theta)=sin(\theta)\qquad thus\\ \quad \\ cos(90^o-20^o)=sin(\square ?)\)

OpenStudy (anonymous):

yeah both are 90-0 =0

OpenStudy (anonymous):

They are both the same. Sin(20) equals roughly .342 and cos (70) roughly equals .342

OpenStudy (jdoe0001):

\(\bf cos(90^o-\theta)=sin(\theta)\qquad thus\\ \quad \\ cos(90^o-20^o)=sin(20^o)\implies cos(70^o)=sin(20^o)\)

OpenStudy (anonymous):

thank you guys so much. i wish i understood this stuff. its just a bunch of letters and numbers to me xD

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

okay so writing it in a complete sentence would be, the sin 20 ad the cos 70 are related because they are equal. cos 90-0 = sin 0 cos 90-20= sin 20 -> cos 70=sin=20 ????

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

Math takes a lot of practice to understand

OpenStudy (anonymous):

i hate it. lol i took algebra 1 in 8th grade and i didnt pass until the beginning of my 10th grade year.